# IntrEx: A Dataset for Modeling Engagement in Educational Conversations

### v1

@@ 1 Introduction
To bridge this gap, we introduce IntrEx, a dataset for modelling interestingness and its expectation in educational conversations, resulting from a comprehensive annotation process, involving over 100 second-language learners, to introduce sequence-level annotations of interestingness that for the first time capture at scale how interestingness and expectations evolve across extended interactions. We introduce two key dimensions of conversational engagement: interestingness, reflecting the perceived level of engagement, and expected interestingness, capturing the anticipated level of engagement. These two scores indicate not only what learners found interesting, but also what they expected to find interesting Murayama (2022). However, assessing interestingness is inherently subjective and prone to variability among annotators. To improve consistency, we thus devised a comparison-based annotation approach inspired by reinforcement learning from human feedback (RLHF) strategies Stiennon et al. (2020); Ouyang et al. (2022). Instead of rating conversational turns in isolation, annotators compare original utterances with automatically rewritten ‘boring’ versions, leading to higher agreement and reduced noise.
Enabled by our new annotations, we conduct two main analyses focused on: (i) the linguistic features that contribute to interestingness, and (ii) the notion of interestingness encoded in LLMs. For linguistic features, we analysed concreteness, comprehensibility, and uptake via a wide range of metrics Hosseini et al. (2022); Lyu et al. (2024), disentangling how each factor influences engagement in teacher-student interactions using linear mixed-effect regressions. We conduct experiments on standard vs. instruction-based fine-tuning and analyze how our dataset improves alignment with human interest judgments. Notably, we find that fine-tuned models with only 77B–88B parameters outperform larger models like GPT-4o OpenAI, (2024), in predicting interestingness in educational dialogues, demonstrating the potential of this new dataset for reward-modelling and for fostering research aimed at improving conversational models for second-language learning. We then conduct a comprehensive investigation of the ability of LLMs to predict interestingness, similarly to human annotators. Recent work (e.g., LaMDA Thoppilan et al. (2022)) has incorporated interestingness as an implicit scoring metric, but no existing dataset provides explicit human-labeled annotations for engagement in educational dialogues.
@@ 3.2 Task Definition
Turn-level Annotations. During the turn-level annotation task, half of the annotators imagined themselves as the teacher in the conversation and rated the interestingness of the student’s responses. The other half imagined themselves as the student in the conversation and rated the interestingness of the teacher’s responses. Our annotation platform (App. F, Fig. 5) displays the conversation page by page, where each page contains a dialogue snippet: a teacher turn and a student turn. Annotators were asked to assign an interestingness score and an expected interestingness score based on the turn from the other party. However, we observed that interestingness scores often remained unchanged across consecutive messages, particularly within the same topic. This led to annotator fatigue, reducing annotation quality and lowering inter-annotator agreement. Additionally, engagement in conversations is not always localized to a single turn, it evolves over longer sequences, making turn-level scores too fine-grained to capture conversational dynamics.
@@ 3.3 Data Collection and Validation
Upon registering for the study, all participants were provided detailed guidelines and informed that the data would be anonymized and made available to researchers. Participants were then directed to our locally hosted annotation platform, which was built using the open-source software doccano [^4] https://github.com/doccano/doccano. Participants were instructed to assume the role of either teacher or student within a given conversation, and to base their ratings on that role, rather than their personal preferences. No participant was assigned both roles within the same conversation.
@@
To conduct this annotation, we leveraged TSCC V2 labels to segment conversations based on teaching content, ensuring that each sequence reflected a cohesive unit of interaction. As described in Section 3.2, annotators assigned interestingness and expected interestingness scores to each sequence. To enhance annotation consistency and agreement, also in this case we incorporated a comparison-based alternative, where annotators evaluated original turns alongside GPT-4o mini-generated rewrites to facilitate more consistent engagement judgments.
@@
where $f_{ij}$ is the frequency with which the first rater gives score $i$ and the second rater gives score $j$, $e_{ij}$ is the expected frequency under the assumption of statistical independence, and $w_{ij}$ is the weight for agreement between scores $i$ and $j$. Using an identity matrix as $W$ would treat both small ($3\rightarrow 4$) and large ($1\rightarrow 4$) differences equally as complete disagreement. Thus, we instead applied a linear weight matrix (Eq. 2), which make disagreement decrease linearly with the distance between categories. It is also necessary to clarify that this choice of weight matrix is part of the evaluation metric and is independent from the data.
@@ 3.4 Dataset Statistics and Demographics
Table 2 presents the self-reported age, gender, first language, and English proficiency (measured using the Common European Framework of Reference, CEFR) of participants in the turn- and sequence-level annotation tasks. In both cases, most annotators identified as B2 or C1 level, reflecting an upper-intermediate to advanced command of English. This distribution was intentional, as it aligns with the proficiency levels of students in the original TSCC dataset (i.e., mostly from B1 to C2), ensuring that annotators could reliably adopt the perspective of the original students in the conversations. We also found that conversations with learners of higher proficiency were rated as more interesting and also that ratings were higher when the annotator proficiency matched the proficiency of the learner in the original conversation. Nonetheless, it is possible that the skew towards annotators with relatively high proficiency may introduce some bias, as annotators not actively engaged in language learning might have perceived the teacher’s instruction as overly simplistic. Additionally, the recruitment platform, Prolific, which is primarily used in Europe and North America, naturally favors users with moderate to advanced English proficiency.
@@ 4 Experiments
In this section, we leverage the new IntrEx dataset, to conduct an experimental assessment of engagement in educational dialogue, focusing on two complementary goals: (i) evaluating the extent to which LLMs align with human judgments of interestingness, and (ii) identifying the linguistic features that contribute to perceived engagement.
@@ 4.2 Linguistic Predictors of Human Interest
  • Coleman-Liau index is depends on the average number of letters per 100 words and the average number of sentences per 100 words.
@@
  • Smog index uses the number of sentences and polysyllables to determine years of education needed to understand text.
@@
  • Student-uptake-teacher assess the extent to which the next speaker built upon the previous speaker’s contribution for two consecutive turns. This metric was derived from a BERT model Wang and Demszky (2024a) trained on a dataset of educational conversations Demszky et al. (2023). The score represent student’s uptake of the teacher’s turn.
@@
Comprehensibility. Comprehensibility was assessed via a wide-range of metrics (i.e., GIS, Flesch reading ease, Flesch-Kincaid grade, Coleman-Liau index, Dale-Chall readability, Gunning fog index, Smog index, Automated readability index, and Spache Readability) We first run comparisons between a simple linear regression model and a regression model including both a linear and a quadratic predictor (linear and quadratic predictors were always orthogonal polynomials) for each comprehensibility feature separately. We then included all significant (i.e., $|t|\geq 2$) predictors in a single linear mixed-effect model with random (intercept) effects for conversation and annotation project (i.e., subset of $3$ annotators who completed the same annotation task). This was done twice: with either interestingness or expected interestingness as the outcome variable and we only retained those metrics that were robust predictors in both models.
@@
At both turn- and sequence-level, several readability metrics were positively related to interest, suggesting that conversations that are more suitable for a higher grade tended to be more interesting for our annotators. Note that, at the sequence-level, some readability metrics had instead a negative relation with interest, suggesting the relation between comprehensibility and interest is complex. In addition, but only at the turn-level, we observed an inverted U-shaped relation with turn length, i.e., lexicon count, the GIS score, and gunning fog, a traditional readability metric. Such inverted U-shaped relation is predicted by several computational and psychological theories of interest Oudeyer et al. (2016); Dubey and Griffiths (2020), and suggests that both very simple and very complex messages lower interest levels, compared to messages that challenge the reader “just enough”.
@@
Uptake. Whule concreteness and comprehensibility are widely applicable to differet types of text, the third feature we investigated, uptake, is specific to conversational turns. Qualitative research using conversation analysis has long emphasized the importance of uptake, particularly in educational settings Huth (2011); Walsh (2013). The degree to which a teacher re-uses elements of a student’s turn provides reassurance that the student is being heard and encourages further interaction, thereby enhancing the learning process and boosting student confidence. See et al. (2019) explored whether controlling for response-relatedness (measured via embedding similarity) improves human judgments of interestingness in human-LSTM network conversations. Inspired by this work, we also investigated response-relatedness, calculating several uptake measures: LCS, propTinS, model-based estimates of conversational uptake, and embedding-based cosine similarity between turns.
@@
In both turn-level and sequence-level analyses, we found positive correlations between some uptake measures and interest (LCS and student-uptake-teacher in turn-level analyses, propTinS in sequence-level analyses). However, cosine similarity was a negative predictor in both turn-level and sequence-level analyses, and student-uptake-teacher was a negative predictor in sequence-level analyses. This seemingly contradictory finding could suggest that introducing novel ideas, which lowers similarity between successive turns or sequences, may also be important to promote interest, and that uptake of the teacher by the student may index a different phenomenon when it is computed at the turn vs. the sequence-level.
@@ 5 Conclusion
In this work, we introduced IntrEx, the first dataset that annotates both interestingness and expected interestingness for teacher–student dialogues, extending TSCC with sequence-level labels to track how engagement evolves across a lesson. We collected ratings from more than hundred L2 learners and designed a comparison-based annotation achieving substantially higher reliability at sequence- than turn-level. Small, instruction-tuned LLMs fine-tuned on IntrEx surpassed GPT-4/4o in predicting human interest ratings, demonstrating the relevance and impact of hiqh quality labels for the task. Future work will be able to leverage IntrEx for training reward models for generation, and extend the approach to new learner groups and domains.

### v2

@@ 1 Introduction
To bridge this gap, we introduce IntrEx, a dataset for modelling interestingness and its expectation in educational conversations, resulting from a comprehensive annotation process, involving over 100 second-language learners, to introduce sequence-level annotations of interestingness that, for the first time, capture at scale how interestingness and expectations evolve across extended interactions. We introduce two key dimensions of conversational engagement: interestingness, reflecting the perceived level of engagement, and expected interestingness, capturing the anticipated level of engagement. These two scores indicate not only what learners found interesting, but also what they expected to find interesting Murayama (2022). However, assessing interestingness is inherently subjective and prone to variability among annotators. To improve consistency, we thus devised a comparison-based annotation approach inspired by reinforcement learning from human feedback (RLHF) strategies Stiennon et al. (2020); Ouyang et al. (2022). Instead of rating conversational turns in isolation, annotators compare original utterances with automatically rewritten ‘boring’ versions, leading to higher agreement and reduced noise.
nabled by our new annotations, we conduct two main analyses focused on: (i) the linguistic features that contribute to interestingness, and (ii) the notion of interestingness encoded in LLMs. For linguistic features, we analysed concreteness, comprehensibility, and uptake via a wide range of metrics Hosseini et al. (2022); Lyu et al. (2024), disentangling how each factor influences engagement in teacher-student interactions using linear mixed-effect regressions. We conduct experiments on standard vs. instruction-based fine-tuning and analyze how our dataset improves alignment with human interest judgments. Notably, we find that fine-tuned models with only 77B–88B parameters outperform larger models like GPT-4o OpenAI, (2024), in predicting interestingness in educational dialogues, demonstrating the potential of this new dataset for reward-modelling and for fostering research aimed at improving conversational models for second-language learning. We then conduct a comprehensive investigation of the ability of LLMs to predict interestingness, similarly to human annotators. Recent work (e.g., LaMDA Thoppilan et al. (2022)) has incorporated interestingness as an implicit scoring metric, but no existing dataset provides explicit human-labeled annotations for engagement in educational dialogues.
@@ 3.2 Task Definition
Turn-level Annotations. During the turn-level annotation task, half of the annotators imagined themselves as the teacher in the conversation and rated the interestingness of the student’s responses. The other half imagined themselves as the student in the conversation and rated the interestingness of the teacher’s responses. Our annotation platform (App. F, Fig. 5) displays the conversation page by page, where each page contains a dialogue snippet: a teacher turn and a student turn. Annotators were asked to assign an interestingness score and an expected interestingness score based on the turn from the other party. However, we observed that interestingness scores often remained unchanged across consecutive messages, particularly within the same topic. This led to annotator fatigue, reducing annotation quality and lowering inter-annotator agreement. Additionally, engagement in conversations is not always localized to a single turn; it evolves over longer sequences, making turn-level scores too fine-grained to capture conversational dynamics.
@@ 3.3 Data Collection and Validation
Upon registering for the study, all participants were provided with detailed guidelines and informed that the data would be anonymized and made available to researchers. Participants were then directed to our locally hosted annotation platform, which was built using the open-source software doccano [^4] https://github.com/doccano/doccano. Participants were instructed to assume the role of either teacher or student within a given conversation, and to base their ratings on that role, rather than their personal preferences. No participant was assigned both roles within the same conversation.
@@
To conduct this annotation, we leveraged TSCC V2 labels to segment conversations based on teaching content, ensuring that each sequence reflected a cohesive unit of interaction. As described in Section 3.2, annotators assigned interestingness and expected interestingness scores to each sequence. To enhance annotation consistency and agreement, also in this case we also incorporated a comparison-based alternative, where annotators evaluated original turns alongside GPT-3.5-generated rewrites to facilitate more consistent engagement judgments.
@@
where $f_{ij}$ is the frequency with which the first rater gives score $i$ and the second rater gives score $j$, $e_{ij}$ is the expected frequency under the assumption of statistical independence, and $w_{ij}$ is the weight for agreement between scores $i$ and $j$. Using an identity matrix as $W$ would treat both small ($3\rightarrow 4$) and large ($1\rightarrow 4$) differences equally as complete disagreement. Thus, we instead applied a linear weight matrix (Eq. 2), which makes disagreement decrease linearly with the distance between categories. It is also necessary to clarify that this choice of weight matrix is part of the evaluation metric and is independent from the data.
@@ 3.4 Dataset Statistics and Demographics
Table 2 presents the self-reported age, gender, first language, and English proficiency (measured using the Common European Framework of Reference, CEFR) of participants in the turn- and sequence-level annotation tasks. In both cases, most annotators identified as B2 or C1 level, reflecting an upper-intermediate to advanced command of English. This distribution was intentional, as it aligns with the proficiency levels of students in the original TSCC dataset (i.e., mostly from B1 to C2), ensuring that annotators could reliably adopt the perspective of the original students in the conversations. We also found that conversations with learners of higher proficiency were rated as more interesting, and also that ratings were higher when the annotator’s proficiency matched the proficiency of the learner in the original conversation. Nonetheless, it is possible that the skew towards annotators with relatively high proficiency may introduce some bias, as annotators not actively engaged in language learning might have perceived the teacher’s instruction as overly simplistic. Additionally, the recruitment platform, Prolific, which is primarily used in Europe and North America, naturally favors users with moderate to advanced English proficiency.
@@ 4 Experiments
In this section, we leverage the new IntrEx dataset to conduct an experimental assessment of engagement in educational dialogue, focusing on two complementary goals: (i) evaluating the extent to which LLMs align with human judgments of interestingness, and (ii) identifying the linguistic features that contribute to perceived engagement.
@@ 4.2 Linguistic Predictors of Human Interest
  • Coleman-Liau index depends on the average number of letters per 100 words and the average number of sentences per 100 words.
@@
  • Smog index uses the number of sentences and polysyllables to determine the years of education needed to understand text.
@@
  • Student-uptake-teacher assess the extent to which the next speaker built upon the previous speaker’s contribution for two consecutive turns. This metric was derived from a BERT model Wang and Demszky (2024a) trained on a dataset of educational conversations Demszky et al. (2023). The score represents the student’s uptake of the teacher’s turn.
@@
Comprehensibility. Comprehensibility was assessed via a wide-range of metrics (i.e., GIS, Flesch reading ease, Flesch-Kincaid grade, Coleman-Liau index, Dale-Chall readability, Gunning fog index, Smog index, Automated readability index, and Spache Readability) We first run comparisons between a simple linear regression model and a regression model including both a linear and a quadratic predictor (linear and quadratic predictors were always orthogonal polynomials) for each comprehensibility feature separately. We then included all significant (i.e., $|t|\geq 2$) predictors in a single linear mixed-effect model with random (intercept) effects for conversation and annotation project (i.e., subset of $3$ annotators who completed the same annotation task). This was done twice: with either interestingness or expected interestingness as the outcome variable, and we only retained those metrics that were robust predictors in both models.
@@
At both turn- and sequence-level, several readability metrics were positively related to interest, suggesting that conversations that are more suitable for a higher grade tended to be more interesting for our annotators. Note that, at the sequence-level, some readability metrics had instead a negative relation with interest, suggesting the relation between comprehensibility and interest is complex. In addition, but only at the turn-level, we observed an inverted U-shaped relation with turn length, i.e., lexicon count, the GIS score, and gunning fog, a traditional readability metric. Such an inverted U-shaped relation is predicted by several computational and psychological theories of interest Oudeyer et al. (2016); Dubey and Griffiths (2020), and suggests that both very simple and very complex messages lower interest levels, compared to messages that challenge the reader “just enough”.
@@
Uptake. While concreteness and comprehensibility are widely applicable to different types of text, the third feature we investigated, uptake, is specific to conversational turns. Qualitative research using conversation analysis has long emphasized the importance of uptake, particularly in educational settings Huth (2011); Walsh (2013). The degree to which a teacher reuses elements of a student’s turn provides reassurance that the student is being heard and encourages further interaction, thereby enhancing the learning process and boosting student confidence. See et al. (2019) explored whether controlling for response-relatedness (measured via embedding similarity) improves human judgments of interestingness in human-LSTM network conversations. Inspired by this work, we also investigated response-relatedness, calculating several uptake measures: LCS, propTinS, model-based estimates of conversational uptake, and embedding-based cosine similarity between turns.
@@
In both turn-level and sequence-level analyses, we found positive correlations between some uptake measures and interest (LCS and student-uptake-teacher in turn-level analyses, propTinS in sequence-level analyses). However, cosine similarity was a negative predictor in both turn-level and sequence-level analyses, and student-uptake-teacher was a negative predictor in sequence-level analyses. This seemingly contradictory finding could suggest that introducing novel ideas, which lowers similarity between successive turns or sequences, may also be important to promote interest, and that uptake of the teacher by the student may index a different phenomenon when it is computed at the turn-level vs. the sequence-level.
@@ 5 Conclusion
We introduced IntrEx, the first dataset that annotates both interestingness and expected interestingness for teacher–student dialogues, extending TSCC with sequence-level labels to track how engagement evolves across a lesson. We collected ratings from more than a hundred L2 learners and designed a comparison-based annotation, achieving substantially higher reliability at the sequence-level than the turn-level. Small, instruction-tuned LLMs fine-tuned on IntrEx surpassed GPT-4/4o in predicting human interest ratings, demonstrating the relevance and impact of hiqh-quality labels for the task. Future work will be able to leverage IntrEx for training reward models for generation, and extend the approach to new learner groups and domains.